Please help! I haven't the foggiest idea how to do an example problem with this formula. I look at it and freeze. Can someone help me? (This is a homework assn.)
If there are “k” possible outcomes for a phenomenon and each is equally likely, then each individual outcome has a probability 1/k. The probability of any event A is:

P(A) = count of Outcomes in A / count of outcomes in S
= count of outcomes in / k

Answer by stanbon(75887)

Please help! I haven't the foggiest idea how to do an example problem with this formula. I look at it and freeze. Can someone help me? (This is a homework assn.)
If there are “k” possible outcomes for a phenomenon and each is equally likely, then each individual outcome has a probability 1/k. The probability of any event A is:
P(A) = count of Outcomes in A / count of outcomes in S
= count of outcomes in / k
---------------------------------
They are trying to defind the probability of an event.
Say you flip a coin.
# of heads is one
# of tails is one
------------------------
The are two possible outcomes.
P(head) = 1/2
P(tail) = 1/2
In this case k = 2
------------------------------
Say you have a die with faces 1,2,3,4,5,6
Here k = 6
There is only one way to get each of the results.
P(1) = 1/6
P(2) = 1/6
P(3) = 1/6
etc.
------------------------------
Say you have a family of three children.
The pattern could be bbb,bgg,gbg,ggb,bbg,bgb,gbb,ggg
So there are 8 possible results.
The probability of having 2 girls in a family of
three is 3/8
Here k = 8
===============================
P(success) = [# of ways to succeed]/[# of possible outcomes]
===============================
Hope this helps.
Cheers,
Stan H.